17 March 2014

Routine PEARL operations

The temperature today in Eureka was -45°C with ice crystals. Outside the PEARL Ridge Lab, it was -22°C. There were a cloudy in the sky, but conditions were mostly sunny. This was the Intensive Phase team's last regular day. It was a successful, uneventful, routine day of running PEARL measurements. Tomorrow, we'll be at the lab, but will also be preparing for our departure the following morning.

In the PEARL Ridge Lab's IR lab, Joseph and Dan took 75 NIR and 64 MIR measurements with the Bruker. Dan took 60 measurements with PARIS.

Over in the UV-Vis Lab, the PEARL-GBS is working properly again. Yesterday's issue was caused by an incorrect positioning of the filter wheel turret. It had been moved unintentionally. The instrument, suntracker, and computer were restarted after the shutter was removed. Everything is fine now with the PEARL-GBS. The UT-GBS and SAOZ instruments continue to run measurements.

Today was the 11th day of continuous CRL measurements at 0PAL. Due to internet difficulties yesterday night the CRL team lost few hours of data. Both channels otherwise took measurements as usual.

Mid-afternoon, the Eureka Weather Station let the PEARL team know that wolves had been spotted near Eureka. Some of the station staff had gone to look. On our way back from the lab, we were fortunate enough to see a wolf as well!

- Dan
Back to Daily Reports
Back to Home